OpenStudy (anonymous):

Jeff is responsible for buying supplies of food and medication for the dogs and cats at a local shelter. The food and medication for a dog cost twice as much as for a cat. He needs to provide supplies for 164 cats and 24 dogs, and his budget is $4,240. How much can he spend on each dog for food and medication

OpenStudy (anonymous):

@Kemist let d = the cost of food and medication for one dog c = the cost of food medication for one cat the food and medication for a dog cost twice as much as for a cat d = 2c divide both sides by 2 c = d/2 he needs to provide supplies for 164 cats and 24 dogs, and his budget is $4,240 24d + 164c <= 4240 plug c = d/2 into the last equation 24d + 164(d/2) <= 4240 by solving the inequality we get d <= $40 am i right?

OpenStudy (anonymous):

what is ur name in real life

OpenStudy (anonymous):

whos

OpenStudy (anonymous):

Correcto

OpenStudy (anonymous):

kemist

OpenStudy (anonymous):

Me? Raven. :3 Nice to meet you.

OpenStudy (anonymous):

boy? girl?

OpenStudy (anonymous):

i thought kemist was a boy :0

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

I am :(

OpenStudy (anonymous):

@Kemist you jus chill like that

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

Just have a girl name :) But I am a guy.

OpenStudy (anonymous):

lol :))

OpenStudy (anonymous):

oh my bday is 12-16-97 yalls?

OpenStudy (anonymous):

My parents were fascinated with the crow so they named me Raven. :(

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

ur funny

OpenStudy (anonymous):

6/6/98

OpenStudy (anonymous):

12-16-1996

OpenStudy (anonymous):

woah that creppy we have the same bdays

OpenStudy (anonymous):

except me im lame ._.

OpenStudy (anonymous):

no ur not

OpenStudy (anonymous):

in 2006 i was the devils child i swear XD 6/6/06 :0

OpenStudy (anonymous):

what makes u think that?

OpenStudy (anonymous):

6/6/6?? lol

OpenStudy (anonymous):

OMG!!!!! i just looked it up and omg thats creppy but ur the devil's baby

OpenStudy (anonymous):

Did you know Santa is Satan in disguise? o_O

OpenStudy (anonymous):

ur NOT the devils baby

OpenStudy (anonymous):

how u know that?

OpenStudy (anonymous):

and what makes u think tht